# Some Super-approximation Rates of ReLU Neural Networks for Korobov Functions

### v1

@@ 4.2.1 The case $m=2$
We first need to construct the desired function $\psi_{\bm{k}}$ to efficiently appropriate the interpolation function
@@
Step 1 Construct the desired function $\psi_{\bm{k}}$ to efficiently appropriate the interpolation function $\Pi_{n}f$.
To this end, we fisrt construct a function $\psi_{\bm{k}}^{\bm{l}}$ for
@@
It has been proved in Theorem 1 that there exists a function
@@
where $|\widetilde{v}_{\bm{l},\bm{i}}-v_{\bm{l},\bm{i}}|\leq 2W^{-2s}(2L)^{-2s}$ and $|\widetilde{v}_{\bm{l},\bm{i}}|\leq 2$ for any $\bm{i}\in\mathcal{I}_{\bm{l}}$. By Lemma 6 with $c=2$, there exists a function $\phi_{2}\in(15(W+1)+2d-1,28d^{2}L)$ such that $\|\phi_{2}\|_{W^{1}_{\infty}\left([0,1]^{d}\times[-2,2]\right)}\leq 12\cdot 2^{d+1}$ and
$\left\|\phi_{2}(\bm{x})-x_{1}x_{2}\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu}x_{d+1}\right\|_{W^{1}_{\infty}\left([0,1]^{d}\times[-2,2]\right)}$ (33)
$\leq$ $14\cdot 2^{4d+4}(W+1)^{-14dL}.$
Define $\phi^{\bm{l}}_{\bm{k}}:=\phi_{2}\circ\phi_{1}$. By Lemma 2,
@@
When $s=3$ and utilizing equation (33) along with the the inequality $\|\phi_{\bm{l},\bm{i}}(\bm{x})\|_{W^{1}_{\infty}(\Omega_{\bm{i},\varepsilon}^{\bm{l}})}\leq W^{2}L^{2}$ implies
@@
$\leq$ $\max_{\bm{i}\in\mathcal{I}_{\bm{l}}}\left(\|\phi^{\bm{l}}_{\bm{k}}(\bm{x})-\widetilde{v}_{\bm{l},\bm{i}}\phi_{\bm{l},\bm{i}}(\bm{x})\|_{W^{1}_{\infty}(\Omega_{\bm{i},\varepsilon}^{\bm{l}})}\right.$
$\left.+|\widetilde{v}_{\bm{l},\bm{i}}-v_{\bm{l},\bm{i}}|\|\phi_{\bm{l},\bm{i}}(\bm{x})\|_{W^{1}_{\infty}(\Omega_{\bm{i},\varepsilon}^{\bm{l}})}\right)$
$\leq$ $\max_{\bm{i}\in\mathcal{I}_{\bm{l}}}\left(\max\left(1,\max_{1\leq j\leq d}|\phi_{l_{j},i_{j}}(x_{j})|_{W^{1}_{\infty}(\Omega_{i_{j},\varepsilon}^{l_{j}})}\right)\right)$
$\cdot C_{d}(W+1)^{-14dL}$
@@
Define $\psi_{\bm{k}}=\sum_{|\bm{l}|_{1}\leq n+d-1}\phi^{\bm{l}}_{\bm{k}}$. According to Lemma 4, we have that
@@
$\|\psi_{\bm{k}}-\Pi_{n}f\|_{W^{1}_{\infty}(\Omega_{\bm{k}})}$ $\leq\sum_{|\bm{l}|_{1}\leq n+d-1}\|\phi^{\bm{l}}_{\bm{k}}(\bm{x})-f_{\bm{l}}(\bm{x})\|_{W^{1}_{\infty}(\Omega_{\bm{k}})}$ (34)
@@
By Lemma 2.1 in Bungartz and Griebel (2004) and the fact that
@@
$\Big{\|}\sum_{\bm{i}\in\mathcal{I}_{\bm{l}}}v_{\bm{l},\bm{i}}\phi_{\bm{l},\bm{i}}(\bm{x})\Big{\|}_{W^{1}_{p}(\Omega)}^{p}$
$=$ $\sum_{\bm{i}\in\mathcal{I}_{\bm{l}}}|v_{\bm{l},\bm{i}}|^{p}\|\phi_{\bm{l},\bm{i}}(\bm{x})\|_{W^{1}_{p}(\Omega_{\bm{l},\bm{i}})}^{p}$
@@
Thus, combining this result with the norm equivalence between $l_{p}$ and $l_{2}$ in $\mathbb{R}^{d}$ and the inequality (proof of (Bungartz and Griebel, 2004, Theorem 3.8))
@@
$\|f-\Pi_{n}f\|_{W^{1}_{p}(\Omega)}$ $\leq C_{p,d}\sum_{|\bm{l}|_{1}>n+d-1}\left(\sum_{j=1}^{d}2^{l_{j}p}\right)^{\frac{1}{p}}\cdot 2^{-2|\bm{l}|_{1}}$ (35)
@@
Combining this result with (7) and (34), we obtain
@@ 4.2.2 The case $m\geq 3$
We need to first construct the desired function $\psi_{\bm{k}}$ to efficiently appropriate the interpolation function
@@
Step 1 Construct the desired function $\psi_{\bm{k}}$ to efficiently appropriate the interpolation function $\Pi_{n}^{m}f$.
@@
It has been demonstrated in the proof of Theorem 1 that there exists a function $\phi_{1}\in\mathcal{NN}(9mdW,14L)$ such that for any $\bm{x}\in\Omega_{\bm{i},\varepsilon}^{\bm{l}}$,
$\phi_{1}(\bm{x})$ $=\left(\psi_{l_{1},i_{1},1}(x_{1}),\ldots,\psi_{l_{1},i_{1},m-1}(x_{1}),\ldots,\right.$
$\quad\left.\psi_{l_{d},i_{d},1}(x_{d}),\ldots,\psi_{l_{d},i_{d},m-1}(x_{d}),\bm{x}^{\top}\right)^{\top},$
where $0\leq\psi_{l_{j},i_{j},k}=2^{-n-d+1}\rho_{l_{j},i_{j},k}\leq 1$ for any $j=1,2,\ldots,d$ and $k=1,2,\ldots,m-1$. By Lemma 3.5 in Hon and Yang (2022) with $k=m-1$, there exists a function $\phi_{2}\in(9(W+1)+m-2,14(m-1)(m-2)dL)$ such that $\|\phi_{2}\|_{W^{1}_{\infty}\left([0,1]^{m-1}\right)}\leq 18$ and
$\left\|\phi_{2}(\bm{x})-x_{1}x_{2}\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu}x_{m-1}\right\|_{W^{1}_{\infty}\left([0,1]^{m-1}\right)}\leq 10m(W+1)^{-7(m-1)dL}.$
Similar to the proof of the case $m\geq 3$ in Theorem 1, we obtain a function $\widetilde{\phi}\in\mathcal{NN}(20mdW,28m^{2}d^{2}L)$ such that for any $\bm{x}\in\Omega_{\bm{i},\varepsilon}^{\bm{l}}$ we have
@@
$\left\|\widetilde{\psi}_{l_{j},i_{j}}(x_{j})-\prod_{k=1}^{m-1}\psi_{l_{j},i_{j},k}(x_{j})\right\|_{W^{1}_{\infty}\left(\Omega_{i_{j},\varepsilon}^{l_{j}}\right)}\leq 10m(W+1)^{-7(m-1)dL},$ (36)
and
$|\widetilde{\psi}_{l_{j},i_{j}}(x_{j})|\leq 2$
for all $j=1,2,\ldots,d$. For $g_{2}\in\mathcal{NN}(4dW+5d,8L+5)$ defined in (18), let $\phi_{3}=g_{2}\circ\widetilde{\phi}$. Then $\phi_{3}\in\mathcal{NN}(20mdW,41m^{2}d^{2}L)$ and for any $\bm{x}\in\Omega_{\bm{i},\varepsilon}^{\bm{l}}$ we have
@@
by Lemma 3, there exists a function
@@
$g_{6}\binom{\bm{z}_{1}}{z_{2}}=\left(\sigma(\bm{z}_{1})+\sigma(-\bm{z}_{1}),\widetilde{\phi}_{4}(z_{2})\right)^{\top},$
for any $\bm{z}_{1}\in\mathbb{R}^{d},z_{2}\in\mathbb{R}$. Using Lemma 2, we have
@@
$|\widetilde{v}^{m}_{\bm{l},\bm{i}}-v^{m}_{\bm{l},\bm{i}}|\leq 2C_{m,d}W^{-2s}(2L)^{-2s},\quad\text{for any}\,\,\bm{i}\in\mathcal{I}_{\bm{l}}.$ (37)
By utilizing Lemma 6 with $c=3C_{m,d}\geq 3m$, we obtain a function $\phi_{6}\in(9(W+1)+2d+1,28md^{2}L)$ such that $\|\phi_{6}\|_{W^{1}_{\infty}\left(\Omega_{c}\right)}\leq C_{m,d}$ and
$\left\|\phi_{6}(\bm{x})-x_{1}x_{2}\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu}x_{d+1}\right\|_{W^{1}_{\infty}\left(\Omega_{c}\right)}\leq C_{m,d}(W+1)^{-14dmL}.$
Let $\phi^{\bm{l}}_{\bm{k}}:=2^{md(n+d-1)}\cdot\phi_{6}\circ\phi_{5}$. Then, by Lemma 2,
@@
$E_{1}$ $\leq C_{m,d}2^{md(n+d-1)}\cdot\max_{\bm{i}\in\mathcal{I}_{\bm{l}}}\max_{1\leq j\leq d}|\widetilde{\psi}_{l_{j},i_{j}}(x_{j})|_{W^{1}_{\infty}(\Omega_{i_{j},\varepsilon}^{l_{j}})}$
$\quad\cdot(W+1)^{-14mdL}$
@@
by taking $s=m+1$ and using inverse estimates and Equation (36) for the second inequality. As for $E_{3}$, by Lemma 7 and (36), we have
@@
$\|\psi_{\bm{k}}-\Pi_{n}^{m}f\|_{W^{1}_{\infty}(\Omega_{\bm{k}})}$ $\leq\sum_{|\bm{l}|_{1}\leq n+d-1}\|\phi^{\bm{l}}_{\bm{k}}(\bm{x})-f_{\bm{l}}(\bm{x})\|_{W^{1}_{\infty}(\Omega_{\bm{k}})}$ (38)
@@
in Bungartz and Griebel (2004) and the facts (see (Bungartz and Griebel, 2004, Lemma 4.6) and Li and Zhang (2025))
@@
$\Big{\|}\sum_{\bm{i}\in\mathcal{I}_{\bm{l}}}v_{\bm{l},\bm{i}}^{m}\phi^{m}_{\bm{l},\bm{i}}(\bm{x})\Big{\|}_{W^{1}_{p}(\Omega)}^{p}$
$=$ $\sum_{\bm{i}\in\mathcal{I}_{\bm{l}}}|v_{\bm{l},\bm{i}}^{m}|^{p}\|\phi^{m}_{\bm{l},\bm{i}}(\bm{x})\|_{W^{1}_{p}(\Omega_{\bm{l},\bm{i}})}^{p}$
@@
where the first inequality is due to the well-known inverse estimates (see (Brenner and Scott, 1994, Lemma 4.5.3) for instance). Thus, similar to (35), we have
@@
Therefore, combining this result with (9), (38), we have
@@
Remark 1.
As indicated in Corollary 2 in Yang et al. (2023b), for the approximation of Sobolev functions $f\in W^{m}_{\infty}(\Omega)$ in the $W^{n}_{\infty}$ norm with $n\geq 2$ using ReLU-$\text{ReLU}^{2}$ DNN (where the activation function applied to each neuron can be chosen as ReLU or its square) with width $O(W\log W)$ and depth $O(L\log L)$, the superconvergence rates $O((WL)^{-2(m-n)/d})$ can be achieved by employing a smoother unit decomposition similar to $g_{\bm{m}}$, i.e., utilizing piecewise polynomials of degree $m-1$. This methodology can actually be used to obtain the approximation of higher order Korobov functions using ReLU-$\text{ReLU}^{2}$ DNNs in the norm of $W^{n,p}$ with $n\geq 2$, with the help of the error estimates of sparse grid interpolation in this norm. This error can actually be computed using the inverse estimates akin to the process of calculating the $W^{1}_{p}$ error.
Remark 2.
Recently, Floor-ReLU networks (i.e., the activation function applied to each neuron can be chosen as ReLU or Floor function) with super approximation power have been introduced in Shen et al. (2021a) to efficiently approximate Hölder functions on $\Omega$. By observing that the proof process in (Shen et al., 2021a, Proposition 3.2) have provided that for $K=W^{L}$ and $\xi_{i}\in[0,1]$ for $i=1,2,\ldots,K$, there exists a function $\psi\in\mathcal{NN}(2W^{2}+5W,L(7L-1)+2)$ such that
$|\psi(i)-\xi_{i}|\leq 2^{-WL},$
we can provide the results for approximating Korobov functions on $\Omega$ by the Floor-ReLU DNN. Similar to ReLU DNN approximation in the norm of $W^{1}_{p}$, establishing a series of results about approximating polynomial by Floor-ReLU DNN in the $W^{1}_{p}$ norm will lead to the approximation results for higher order Korobov functions using Floor-ReLU DNNs, which will be left for future research.
Remark 3.
Recently, the authors of Liu et al. (2024a) have shown that ResNet (He et al. (2016); Lin and Jegelka (2018); Oono and Suzuki (2019); Li et al. (2022)) with constant width and depth $O(L)$ can approximate Lipschitz continuous function $f\in C([0,1]^{d})$ with the optimal approximation rates $O(\omega_{f}(L^{-2/d})$, where $\omega_{f}$ is the modulus of continuity of the function $f$. Additionally, the versions of ResNet similar to Lemma 1, Lemma 3, and Lemma 6 for $L_{\infty}$ norm are presented. By combining these results with the representation of basis functions by ReLU networks as demonstrated in the proof of Theorem 1, one can derive the approximation rates for Korobov spaces using ResNet through analogous methodologies. For brevity, we will not delve into the proof here. Moreover, we can obtain approximation results of ResNet for Korobov functions in the $W^{1}_{p}$ norm, resembling those outlined in Theorem 2, by formulating the $W^{1}_{p}$ ResNet counterparts of Lemma 1, 3, and 6.

### v2

@@ 4.2.1 The case $m=2$
We first need to construct the desired function $\psi_{\bm{k}}$ to efficiently approximate the interpolation function
@@
Step 1 Construct the desired function $\psi_{\bm{k}}$ to efficiently approximate the interpolation function $\Pi_{n}f$.
To this end, we first construct a function $\psi_{\bm{k}}^{\bm{l}}$ for
@@
It has been proved in Theorem 1.2 that there exists a function
@@
where $|\widetilde{v}_{\bm{l},\bm{i}}-v_{\bm{l},\bm{i}}|\leq 2W^{-2s}(2L)^{-2s}$ and $|\widetilde{v}_{\bm{l},\bm{i}}|\leq 2$ for any $\bm{i}\in\mathcal{I}_{\bm{l}}$. By Lemma 2.7 with $c=2$, there exists a function $\phi_{2}\in(15(W+1)+2d-1,28d^{2}L)$ such that $\|\phi_{2}\|_{W^{1}_{\infty}\left([0,1]^{d}\times[-2,2]\right)}\leq 12\cdot 2^{d+1}$ and
$\left\|\phi_{2}(\bm{x})-x_{1}x_{2}\cdots x_{d+1}\right\|_{W^{1}_{\infty}\left([0,1]^{d}\times[-2,2]\right)}\leq 14\cdot 2^{4d+4}(W+1)^{-14dL}.$ (4.33)
Define $\phi^{\bm{l}}_{\bm{k}}:=\phi_{2}\circ\phi_{1}$. By Lemma 2.2,
@@
When $s=3$ and utilizing equation (4.33) along with the the inequality $\|\phi_{\bm{l},\bm{i}}(\bm{x})\|_{W^{1}_{\infty}(\Omega_{\bm{i},\varepsilon}^{\bm{l}})}\leq W^{2}L^{2}$ implies
@@
$\leq$ $\max_{\bm{i}\in\mathcal{I}_{\bm{l}}}\left(\|\phi^{\bm{l}}_{\bm{k}}(\bm{x})-\widetilde{v}_{\bm{l},\bm{i}}\phi_{\bm{l},\bm{i}}(\bm{x})\|_{W^{1}_{\infty}(\Omega_{\bm{i},\varepsilon}^{\bm{l}})}+|\widetilde{v}_{\bm{l},\bm{i}}-v_{\bm{l},\bm{i}}|\|\phi_{\bm{l},\bm{i}}(\bm{x})\|_{W^{1}_{\infty}(\Omega_{\bm{i},\varepsilon}^{\bm{l}})}\right)$
$\leq$ $\max_{\bm{i}\in\mathcal{I}_{\bm{l}}}\left(\max\left(1,\max_{1\leq j\leq d}|\phi_{l_{j},i_{j}}(x_{j})|_{W^{1}_{\infty}(\Omega_{i_{j},\varepsilon}^{l_{j}})}\right)\right)\cdot C_{d}(W+1)^{-14dL}$
@@
Define $\psi_{\bm{k}}=\sum_{|\bm{l}|_{1}\leq n+d-1}\phi^{\bm{l}}_{\bm{k}}$. According to Lemma 2.4, we have that
@@
$\|\psi_{\bm{k}}-\Pi_{n}f\|_{W^{1}_{\infty}(\Omega_{\bm{k}})}$ $\leq\sum_{|\bm{l}|_{1}\leq n+d-1}\|\phi^{\bm{l}}_{\bm{k}}(\bm{x})-f_{\bm{l}}(\bm{x})\|_{W^{1}_{\infty}(\Omega_{\bm{k}})}$ (4.34)
@@
By Lemma 2.1 in [9] and the fact that
@@
$\Big\|\sum_{\bm{i}\in\mathcal{I}_{\bm{l}}}v_{\bm{l},\bm{i}}\phi_{\bm{l},\bm{i}}(\bm{x})\Big\|_{W^{1}_{p}(\Omega)}^{p}$ $=\sum_{\bm{i}\in\mathcal{I}_{\bm{l}}}|v_{\bm{l},\bm{i}}|^{p}\|\phi_{\bm{l},\bm{i}}(\bm{x})\|_{W^{1}_{p}(\Omega_{\bm{l},\bm{i}})}^{p}$
@@
Thus, combining this result with the norm equivalence between $l_{p}$ and $l_{2}$ in $\mathbb{R}^{d}$ and the inequality (proof of [9, Theorem 3.8])
@@
$\|f-\Pi_{n}f\|_{W^{1}_{p}(\Omega)}$ $\leq C_{p,d}\sum_{|\bm{l}|_{1}>n+d-1}\left(\sum_{j=1}^{d}2^{l_{j}p}\right)^{\frac{1}{p}}\cdot 2^{-2|\bm{l}|_{1}}$ (4.35)
@@
Combining this result with (2.7) and (4.34), we obtain
@@ 4.2.2 The case $m\geq 3$
We need to first construct the desired function $\psi_{\bm{k}}$ to efficiently approximate the interpolation function
@@
Step 1 Construct the desired function $\psi_{\bm{k}}$ to efficiently approximate the interpolation function $\Pi_{n}^{m}f$.
@@
It has been demonstrated in the proof of Theorem 1.2 that there exists a function $\phi_{1}\in\mathcal{NN}(9mdW,14L)$ such that for any $\bm{x}\in\Omega_{\bm{i},\varepsilon}^{\bm{l}}$,
$\phi_{1}(\bm{x})$ $=\left(\psi_{l_{1},i_{1},1}(x_{1}),\ldots,\psi_{l_{1},i_{1},m-1}(x_{1}),\ldots,\psi_{l_{d},i_{d},1}(x_{d}),\ldots,\psi_{l_{d},i_{d},m-1}(x_{d}),\bm{x}^{\top}\right)^{\top},$
where $0\leq\psi_{l_{j},i_{j},k}=2^{-n-d+1}\rho_{l_{j},i_{j},k}\leq 1$ for any $j=1,2,\ldots,d$ and $k=1,2,\ldots,m-1$. By Lemma 3.5 in [17] with k=m1k=m-1, there exists a function $\phi_{2}\in(9(W+1)+m-2,14(m-1)(m-2)dL)$ such that $\|\phi_{2}\|_{W^{1}_{\infty}\left([0,1]^{m-1}\right)}\leq 18$ and
$\left\|\phi_{2}(\bm{x})-x_{1}x_{2}\cdots x_{m-1}\right\|_{W^{1}_{\infty}\left([0,1]^{m-1}\right)}\leq 10m(W+1)^{-7(m-1)dL}.$
Similar to the proof of the case $m\geq 3$ in Theorem 1.2, we obtain a function
$\widetilde{\phi}\in\mathcal{NN}(20mdW,28m^{2}d^{2}L)$
such that for any $\bm{x}\in\Omega_{\bm{i},\varepsilon}^{\bm{l}}$ we have
@@
[…]
for all $j=1,2,\ldots,d$. For $g_{2}\in\mathcal{NN}(4dW+5d,8L+5)$ defined in (3.18), let $\phi_{3}=g_{2}\circ\widetilde{\phi}$. Then $\phi_{3}\in\mathcal{NN}(20mdW,41m^{2}d^{2}L)$ and for any $\bm{x}\in\Omega_{\bm{i},\varepsilon}^{\bm{l}}$ we have
@@
by Lemma 2.3, there exists a function
@@
$g_{6}\binom{\bm{z}_{1}}{z_{2}}=\left(\sigma(\bm{z}_{1})-\sigma(-\bm{z}_{1}),\widetilde{\phi}_{4}(z_{2})\right)^{\top},$
for any $\bm{z}_{1}\in\mathbb{R}^{d},z_{2}\in\mathbb{R}$. Using Lemma 2.2, we have
@@
$|\widetilde{v}^{m}_{\bm{l},\bm{i}}-v^{m}_{\bm{l},\bm{i}}|\leq 2C_{m,d}W^{-2s}(2L)^{-2s},\quad\text{for any}\,\,\bm{i}\in\mathcal{I}_{\bm{l}}.$ (4.37)
By utilizing Lemma 2.7 with $c=3C_{m,d}\geq 3m$, we obtain a function $\phi_{6}\in(9(W+1)+2d+1,28md^{2}L)$ such that $\|\phi_{6}\|_{W^{1}_{\infty}\left(\Omega_{c}\right)}\leq C_{m,d}$ and
$\left\|\phi_{6}(\bm{x})-x_{1}x_{2}\cdots x_{d+1}\right\|_{W^{1}_{\infty}\left(\Omega_{c}\right)}\leq C_{m,d}(W+1)^{-14dmL}.$
Let $\phi^{\bm{l}}_{\bm{k}}:=2^{md(n+d-1)}\cdot\phi_{6}\circ\phi_{5}$. Then, by Lemma 2.2,
@@
$E_{1}$ $\leq C_{m,d}2^{md(n+d-1)}\cdot\max_{\bm{i}\in\mathcal{I}_{\bm{l}}}\max_{1\leq j\leq d}|\widetilde{\psi}_{l_{j},i_{j}}(x_{j})|_{W^{1}_{\infty}(\Omega_{i_{j},\varepsilon}^{l_{j}})}\cdot(W+1)^{-14mdL}$
@@
by taking $s=m+1$ and using inverse estimates and Equation (4.36) for the second inequality. As for $E_{3}$, by Lemma 2.9 and (4.36), we have
@@
$\|\psi_{\bm{k}}-\Pi_{n}^{m}f\|_{W^{1}_{\infty}(\Omega_{\bm{k}})}$ $\leq\sum_{|\bm{l}|_{1}\leq n+d-1}\|\phi^{\bm{l}}_{\bm{k}}(\bm{x})-f_{\bm{l}}^{m}(\bm{x})\|_{W^{1}_{\infty}(\Omega_{\bm{k}})}$ (4.38)
@@
in [9] and the facts (see [9, Lemma 4.6] and [22])
@@
$\Big\|\sum_{\bm{i}\in\mathcal{I}_{\bm{l}}}v_{\bm{l},\bm{i}}^{m}\phi^{m}_{\bm{l},\bm{i}}(\bm{x})\Big\|_{W^{1}_{p}(\Omega)}^{p}=$ $\sum_{\bm{i}\in\mathcal{I}_{\bm{l}}}|v_{\bm{l},\bm{i}}^{m}|^{p}\|\phi^{m}_{\bm{l},\bm{i}}(\bm{x})\|_{W^{1}_{p}(\Omega_{\bm{l},\bm{i}})}^{p}$
@@
where the first inequality is due to the well-known inverse estimates (see [7, Lemma 4.5.3] for instance). Thus, similar to (4.35), we have
@@
Therefore, combining this result with (2.9), (4.38), we have
@@
Remark 4.7.
For $f\in X^{m}_{\infty}(\Omega)$, it is worth noting that the error of approximating sparse grid interpolation function using ReLU DNNs in the proof of Proposition 4.5 and the used lemmas are all valid for the norm $W^{1}_{\infty}$. Therefore, once the sparse grid interpolation error has been established in the $W^{1}_{\infty}$ norm, the corresponding approximation error for ReLU DNNs applied to $f\in X^{m}_{\infty}(\Omega)$ can be derived under the $W^{1}_{\infty}$ norm.